# Directional Convergence, Benign Overfitting of Gradient Descent in leaky ReLU two-layer Neural Networks

### v1

@@ Appendix A Proof of Directional Convergence
Since all the assumptions of Theorem 4.4 are satisfied, the desired conclusions hold with probability at least $1-2\delta$ under condition (i).
Under condition (ii), the desired conclusions follow from Theorem 4.9. To see this, first note that $\tilde{\varepsilon}_{1}\leq\tfrac{1}{2}$ implies that $\frac{R_{max}^{2}(\boldsymbol{z})}{R_{min}^{2}(\boldsymbol{z})}\leq\frac{1+\frac{\tilde{\varepsilon}_{1}}{3}}{1-\frac{\tilde{\varepsilon}_{1}}{3}}\leq\frac{7}{5}$. So, we have $C_{w}$ in Theorem 4.9 with $C_{w}\leq\tfrac{168e}{5\gamma^{2}}$. So, by setting $\varepsilon_{i}=\tfrac{5\gamma^{2}}{368e},i=1,3$, $C_{w}\varepsilon_{i}\leq\tfrac{1}{2}$ is ensured. Moreover, by setting $\varepsilon_{2}=\tfrac{1}{2}$, $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$ is ensured.
@@
Since all the assumptions of Theorem 4.9 are satisfied, the desired conclusions hold with with probability at least $1-2\delta$ under condition (ii). ∎
Application of Theorem B.11 on the model (sG) is given below:
Theorem C.6.
Consider model (sG). Suppose $n\geq 4C_{3}^{2}\log(\frac{1}{\delta})$ and one of the following holds:
@@
    $q_{+}q_{-}=0$, $n\geq\frac{16(1+\gamma)^{2}}{\gamma^{4}}$, $n\|\boldsymbol{\mu}\|^{2}<\frac{\gamma}{1+(1+\gamma)/2}{\rm tr}(\Sigma)$,
@@
    $\|\boldsymbol{\mu}\|^{2}$ $\geq\frac{16C_{2}(1+\gamma)}{3\gamma^{2}}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\sqrt{\log(\tfrac{1}{\delta})},$
@@
    $q_{+}q_{-}\neq 0$, $n\geq\frac{16}{q_{\gamma}^{2}}$, $n\|\boldsymbol{\mu}\|^{2}\leq\lambda{\rm tr}(\Sigma)$,
    ${\rm tr}(\Sigma)$ $\geq\frac{4(\frac{3\lambda}{4}+1)^{2}}{q_{\gamma}^{2}}n\sqrt{\log(\tfrac{1}{\delta})}\max\left\{C_{1}\|\Sigma\|_{F},C_{1}\|\Sigma\|\sqrt{n\log(\tfrac{1}{\delta})},C_{2}\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\right\},\text{and}$
    $\|\boldsymbol{\mu}\|^{2}$ $\geq\frac{16C_{2}(\frac{3\lambda}{4}+1)^{2}}{3q_{\gamma}^{2}}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\sqrt{\log(\tfrac{1}{\delta})},$
    for some $\lambda\leq\frac{2q_{\gamma}}{3\gamma(1-\gamma)}$,
@@
    $q_{\gamma}>\gamma$, $n\geq\frac{4}{q_{\gamma}^{2}}$, $n\|\boldsymbol{\mu}\|^{2}\geq\lambda{\rm tr}(\Sigma)$,
    ${\rm tr}(\Sigma)$ $\geq\frac{2\sqrt{2}}{q_{\gamma}\sqrt{\tilde{C}(\gamma,q_{+},\lambda)}}n\|\boldsymbol{\mu}\|\{\log(\tfrac{1}{\delta})\}^{1/4}$
    ×$\qquad\times\sqrt{\max\left\{C_{1}\|\Sigma\|_{F},C_{1}\|\Sigma\|\sqrt{n\log(\tfrac{1}{\delta})},C_{2}\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\right\}},\text{and}$
    ${\rm tr}(\Sigma)$ $\geq\frac{C_{2}}{3q_{\gamma}^{2}\tilde{C}(\gamma,q_{+},\lambda)}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\sqrt{n\log(\tfrac{1}{\delta})}$
    for some $\lambda\geq\tfrac{\gamma}{4}$.
@@
By Lemma C.2 and Corollary C.4, event $\bigcap_{i=1,2,3}\tilde{E}_{i}$ holds with probability at least $1-3\delta$ with
@@
Then, the desired conclusion follows by simply rewriting assumptions of Theorem B.11. ∎
We can now see that Theorem 6.1 and 6.2 follow from Theorem C.5 and C.6 by ensuring both assumptions are satisfied. Specifically, Theorem C.5 under condition (i) and Theorem C.6 under condition (iii) imply Theorem 6.1. Similarly, Theorem C.5 under condition (i) and Theorem C.6 under condition (i) and (ii) imply Theorem 6.2 under condition (ii). Lastly, Theorem C.5 under condition (ii) and Theorem C.6 under condition (i) and (ii) imply Theorem 6.2 under condition (ii)

### v2

@@ Appendix B Proof of Directional Convergence
Since all the assumptions of (i) of Theorem 4.8 are satisfied, the desired conclusions hold with probability at least $1-2\delta$ under condition (i).
Next we show the desired conclusions under condition (ii). To see this, first note that $\tilde{\varepsilon}_{1}\leq\tfrac{1}{2}$ implies that $\frac{R_{max}^{2}(\boldsymbol{z})}{R_{min}^{2}(\boldsymbol{z})}\leq\frac{1+\frac{\tilde{\varepsilon}_{1}}{3}}{1-\frac{\tilde{\varepsilon}_{1}}{3}}\leq\frac{7}{5}$. So, we have $C_{w}$ in Theorem 4.8 with $C_{w}\leq\tfrac{168e}{5\gamma^{2}}$. So, by setting $\varepsilon_{i}=\tfrac{5\gamma^{2}}{368e},i=1,3$, $C_{w}\varepsilon_{i}\leq\tfrac{1}{2}$ is ensured. Moreover, by setting $\varepsilon_{2}=\tfrac{1}{2}$, $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$ is ensured.
@@
Since all the assumptions of (ii) of Theorem 4.8 are satisfied, the desired conclusions hold with with probability at least $1-2\delta$ under condition (ii). ∎
Application of Theorem C.13 on the model (sG) is given below:
Theorem D.6.
Consider model (sG). Suppose $n\geq\max\left\{4C_{3}^{2}\log(\frac{1}{\delta}),N(\gamma,q_{+})\right\}$ and one of the following holds:
@@
    $q_{+}q_{-}=0$, $n\|\boldsymbol{\mu}\|^{2}<\frac{\gamma}{1+(1+\gamma)/2}{\rm tr}(\Sigma)$,
@@
    $\|\boldsymbol{\mu}\|^{2}$ $\geq\tilde{D}^{-1}C_{2}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\sqrt{\log(\tfrac{1}{\delta})},$
@@
    $q_{+}q_{-}\neq 0$, $n\|\boldsymbol{\mu}\|^{2}\leq\frac{q_{\gamma}}{2\gamma(1-\gamma)}{\rm tr}(\Sigma)$,
    ${\rm tr}(\Sigma)$ $\geq 20q_{\gamma}^{-2}n\sqrt{\log(\tfrac{1}{\delta})}\max\left\{C_{1}\|\Sigma\|_{F},C_{1}\|\Sigma\|\sqrt{n\log(\tfrac{1}{\delta})},C_{2}\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\right\},\text{and}$
    $\|\boldsymbol{\mu}\|^{2}$ $\geq\tilde{D}^{-1}C_{2}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\sqrt{\log(\tfrac{1}{\delta})},$
@@
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq\frac{q_{\gamma}}{2\gamma(1-\gamma)}{\rm tr}(\Sigma)$,
    ${\rm tr}(\Sigma)$ $\geq\frac{2\sqrt{10}}{q_{\gamma}\sqrt{q_{\gamma}-\gamma}}n\|\boldsymbol{\mu}\|\{\log(\tfrac{1}{\delta})\}^{1/4}$
    $\qquad\times\sqrt{\max\left\{C_{1}\|\Sigma\|_{F},C_{1}\|\Sigma\|\sqrt{n\log(\tfrac{1}{\delta})},C_{2}\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\right\}}.$
@@
By Lemma D.2 and Corollary D.4, event $\bigcap_{i=1,2,3}\tilde{E}_{i}$ holds with probability at least $1-3\delta$ with
@@
Then, the desired conclusion follows by simply rewriting assumptions of Theorem C.13. ∎
We can now see that Theorem 6.2 follow from Theorem D.5 and D.6 by ensuring both assumptions are satisfied. Specifically, Theorem D.5 under condition (i) and Theorem D.6 under condition (i) and (ii) imply Theorem 6.2 under (a)-(i). Similarly, Theorem D.5 under condition (ii) and Theorem D.6 under condition (i) and (ii) imply Theorem 6.2 under condition (a)-(ii). Lastly, Theorem D.5 under condition (i) and Theorem D.6 under condition (iii) imply Theorem 6.2 under condition (b).
D.2 Proof of Theorem 6.3
For model (PM), we have instead of Lemma D.2
Lemma D.7 (Lemma S3.4, S3.5, Hashimoto et al. [2025]).
Consider model (PM). There exist a constant $C(r,K)$ such that for any $\delta\in(0,\tfrac{1}{2})$, event $\tilde{E}_{1}\cap\tilde{E}_{2}$ holds with probability at least $1-2\delta$, where
$R$ $={\rm tr}(\Sigma),$
$\tilde{\varepsilon}_{1}$ $:=C(r,K)\delta^{-\frac{2}{r}}\max\left\{n^{\frac{2}{r}}p^{\frac{2}{r}-\frac{1}{2}},n^{\frac{4}{r}}\right\}\frac{\|\Sigma\|_{F}}{{\rm tr}(\Sigma)},$
$\tilde{\varepsilon}_{2}$ $:=\sqrt{\frac{n}{\delta}}\frac{\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|}{{\rm tr}(\Sigma)}.$
Similar arguments using D.7 instead of D.2 allows us to establish results analogous to Theorem D.5 and D.6.
Theorem D.8.
Consider model (PM). Let $\delta\in(0,\frac{1}{2})$. Suppose, for sufficiently large constant $C$ which depends on $\delta,\gamma,q_{+},r,K$,
${\rm tr}(\Sigma)\geq C\max\left\{n^{\frac{2}{r}}p^{\frac{2}{r}-\frac{1}{2}}\|\Sigma\|_{F},n^{\frac{4}{r}}\|\Sigma\|_{F},n\frac{\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|^{2}}{\|\boldsymbol{\mu}\|^{2}}\right\}$ (101)
and further suppose one of the following conditions:
  1. (i)
    $\|\boldsymbol{\mu}\|^{2}\geq C\max\big\{n^{\frac{2}{r}}p^{\frac{2}{r}-\frac{1}{2}}\|\Sigma\|_{F},n^{\frac{4}{r}}\|\Sigma\|_{F},\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\big\}$,
  2. (ii)
    ${\rm tr}(\Sigma)\geq C\max\big\{n\|\boldsymbol{\mu}\|^{2},n^{\frac{2}{r}+1}p^{\frac{2}{r}-\frac{1}{2}}\|\Sigma\|_{F},n^{\frac{4}{r}+1}\|\Sigma\|_{F},n^{\frac{3}{2}}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\big\}$,
Then, we have with probability at least $1-2\delta$, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, and converges in the direction given in Theorem 4.8.
Proof.
The theorem follows from Theorem 4.8, Lemma D.1, and Lemma D.7 by repeating the same argument from the proof of Theorem D.5 by replacing Lemma D.2 with Lemma D.7. ∎
Similarly, we obtain a result analogous to Theorem D.6.
Theorem D.9.
Consider model (PM). Let $\delta\in(0,\frac{1}{3})$. Suppose $n\geq\max\left\{4C_{3}^{2}\log(\frac{1}{\delta}),N(\gamma,q_{+})\right\}$ and one of the following holds for sufficiently large constant $C$ which depends on $\delta,\gamma,q_{+},r,K$:
  1. (i)
    ${\rm tr}(\Sigma)\geq C\max\{n\|\boldsymbol{\mu}\|^{2},n^{\frac{2}{r}+\frac{1}{2}}p^{\frac{2}{r}-\frac{1}{2}}\|\Sigma\|_{F},n^{\frac{4}{r}+\frac{1}{2}}\|\Sigma\|_{F},n^{\frac{3}{2}}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\}$ and
    $\|\boldsymbol{\mu}\|^{2}\geq C\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|$,
  2. (ii)
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq\frac{q_{\gamma}}{2\gamma(1-\gamma)}{\rm tr}(\Sigma)$, and
    ${\rm tr}(\Sigma)$ $\geq C\|\boldsymbol{\mu}\|\max\left\{n^{\frac{1}{r}+\frac{3}{4}}p^{\frac{1}{r}-\frac{1}{4}}\|\Sigma\|_{F}^{\frac{1}{2}},n^{\frac{2}{r}+\frac{3}{4}}\|\Sigma\|_{F}^{\frac{1}{2}},n^{\frac{5}{4}}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|^{\frac{1}{2}}\right\}.$
Then, we have with probability at least $1-3\delta$
where $c$ is a constant which depends only on $\gamma$ and $q_{+}$.
Proof.
The theorem follows from Theorem C.13, Lemma D.1, Lemma D.7, and Lemma D.4 by repeating the same argument from the proof of Theorem D.6 by replacing Lemma D.2 with Lemma D.7. ∎
Putting these together with $\Sigma=I_{p}$, we obtain Theorem 6.3. Specifically, Theorem 6.3 under condition (i) follows from Theorem D.8 under (i) and Theorem D.9 under (i), Theorem 6.3 under condition (ii) follows from Theorem D.8 under (i) and Theorem D.9 under (ii), and Theorem 6.3 under condition (iii) follows from Theorem D.8 under (ii) and Theorem D.9 under (i).